OpenStudy (anonymous):

Solve 7(x - 4) + 4x = 71. 11 9 7 5

OpenStudy (shaik0124):

7x-28+4x=71 11x-28+28=71+28 11x=99 x=99/11 x=9

OpenStudy (anonymous):

thx for ur help!!! @shaik0124